OpenStudy (anonymous):

Which of the following numbers is correctly represented by 3 x 10-5?(10 to the fifth power) 1).0.00003 2).30,000 3).0.0003

OpenStudy (sidsiddhartha):

0.00003=3*10-5 0.0003=3*10-4 0.003=3*10-3 like this

OpenStudy (anonymous):

So wait 1 2 or 3?

OpenStudy (anonymous):

@Oreo_Kitten, are you writing an online test and just looking for the correct answers? You needto try to develop an understanding. That way you are learning and not just trying to score well on tests.

OpenStudy (anonymous):

Does it matter, I'm retarded thank you good bye

OpenStudy (anonymous):

The problem is 3x10^-5. So your gonna move the decimal to the left 5 times. So the answer would be 0.00003